OpenStudy (anonymous):

need help trying to set up this problem 6 1/4 gallons = how many fl oz ? this is kind of the setup instructor taught me but cant finish it plan is convert gal to fl oz

OpenStudy (anonymous):

\[6.25 gallons \times \frac{ 128 ounces }{ gallons }\] Two things. Notice how "gallons" cancels, and 2) you can multiply your original question amount by this second factor, because it is in essence just multiplying by "1" since it is an identity.

OpenStudy (anonymous):

ik you convert the mixed number into an improper fraction which would make it 25/4

OpenStudy (anonymous):

You can make that conversion if you are asked to do so by your teacher, but it isn't necessary unless you are asked to do that.

OpenStudy (anonymous):

and I remember canceling out gallons

OpenStudy (anonymous):

thats what were supposed to do I wish I could show the setup she showed us but for some reason I cant figure this one out

OpenStudy (anonymous):

So, you are left with 25/4 times 128 ounces. You can simplify this by factoring the 128 to get 25/4 times (4 x 32) ounces = 25 times 32 ounces and then just multiply 25 by 32.

OpenStudy (anonymous):

oh ok thank you